### v1

# On the self-consistency of compact objects in Lorentz-violating gravity theories

Abstract
Self-consistent solutions in Lorentz-violating gravity theories require the simultaneous satisfaction of: (i) the corresponding Einstein field equations, (ii) the matter field equations, and (iii) the Lorentz-violating field equations. In vacuum states, the dynamics of Lorentz-violating tensor fields may reduce to geometric constraints, potentially precluding entire classes of compact objects. These constraints are crucial for ensuring physical consistency in Lorentz-violating frameworks, as they eliminate metric families incompatible with the anisotropies induced by spontaneous Lorentz symmetry breaking. We investigate the criteria governing the emergence of these geometric constraints and analyze their consequences. Our analysis establishes a consistency framework for evaluating compact objects in these theories, demonstrating that several previously reported solutions in Lorentz-violating gravity models are physically inadmissible.
Introduction. General relativity (GR) remains the most successful theory for describing gravitational interaction at classical scales, grounded in the symmetries of diffeomorphism and local Lorentz invariance. However, the unification of GR with Quantum Mechanics still represents one of the greatest unsolved challenges in theoretical physics. This motivates the plausible conjecture that modifications to the foundations of GR may pave the way toward a unified theory. The typical energy scale at which such a unified framework should becomes relevant is at the so-called Planck scale ($M_{\text{Pl}}=\sqrt{\hbar c/G}\approx 1.22\times 10^{19}\text{ GeV}/c^{2}$), far exceeding any observational regime accessible today. Nevertheless, theoretical extensions of GR predict the existence of low-energy “relic signatures” [1, 2].
These signatures, which might arise from quantum gravitational effects, could manifest as subtle deviations either in experiments of the Standard Model or GR. In this context, the possibility of Lorentz symmetry violation in an underlying theory—such as string theory [3, 4]—becomes particularly relevant, especially in the search for observational signatures of new physics. These questions can be investigated independently of specific models through effective field theory approaches [5]. Thus, experimental tests of Lorentz violation in gravity not only validate extensions of GR but may also reveal evidence of physics beyond the current paradigm [6, 7].
A promising approach to investigate Lorentz violation in pure gravity involves analyzing anisotropy effects induced in compact objects by tensor fields that acquire a preferred spacetime direction—a phenomenon arising from spontaneous Lorentz symmetry breaking [8]. In this mechanism, a potential term drives the dynamic evolution of the tensor field, leading to a nonzero vacuum expectation value (VEV). While being spontaneous, this violation preserves the Bianchi identities, despite implying both local Lorentz and diffeomorphism violation [5]. This compatibility with the geometric structure of GR makes spontaneous Lorentz symmetry breaking a natural candidate for exploring violation effects in strong gravity regimes, such as those found in neutron stars or black holes.
The aim of this letter is to stress out a point, sometimes missing in the discussion on compact objects in Lorentz-violating gravity models, namely the self-consistency of compact objects in these frameworks. Crucially, self-consistent solutions of Lorentz-violating scenarios must simultaneously satisfy: (i) the corresponding Einstein field equations, (ii) the matter field equations, and (iii) the Lorentz-violating field equations. Remarkably, under certain conditions, the latter in a vacuum state can reduce to geometric constraints governing the metrics of compact objects. These constraints impose stringent requirements that rule out the emergence of families of compact objects and invalidate some previously reported solutions of Lorentz-violating gravity models.
In the following, we analyze the emergence and implications of geometric constraints induced by spontaneous Lorentz violation in two prominent frameworks: (i) via a vector field (the bumblebee field) [5, 8], and (ii) through an antisymmetric rank-2 tensor field (sometimes known as the Kalb-Ramond field) [9]. The non-zero VEV of these fields introduces spacetime anisotropies, leading to small deviations in compact object solutions—including black holes [10, 11, 12, 13], neutron stars [14], and wormholes [15]—compared to standard GR predictions. Our analysis plays a crucial role in the search for viable compact objects within Lorentz-violating gravity, as it allows us to probe various families of metric functions (including those mentioned above) and determine which are permitted by the anisotropies induced by Lorentz-violating fields.
Lorentz-violating framework. The action that encodes Lorentz-violating effects from a vector field (bumblebee field) [5] and from an antisymmetric rank-2 tensor [9] in Riemann-spacetime limit can be written as
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa}(R-2\Lambda)+\mathcal{L}_{\mathfrak{B}}+\mathcal{L}_{B}\right]+S_{m},$ (1)
where $S_{m}=S_{m}[g_{\mu\nu},\psi_{m}]$ is the matter action, $R$ is the Ricci scalar and $\Lambda$ is the cosmological constant. $\mathcal{L}_{\mathfrak{B}}$ and $\mathcal{L}_{B}$ are, respectively, the Lagrangians of the bumblebee and of the antisymmetric rank-2 tensor, given by
$\mathcal{L}_{\mathfrak{B}}=$ $-\frac{1}{4}\mathfrak{B}_{\mu\nu}\mathfrak{B}^{\mu\nu}-\tilde{V}(Y)$
$+\frac{\tilde{\xi}_{1}}{2\kappa}\mathfrak{B}_{\mu}\mathfrak{B}^{\mu}R+\frac{\tilde{\xi}_{2}}{2\kappa}\mathfrak{B}^{\mu}\mathfrak{B}^{\nu}R_{\mu\nu},$ (2)
$\mathcal{L}_{B}=$ $-\frac{1}{12}H_{\lambda\mu\nu}H^{\lambda\mu\nu}-V(X)+\frac{\xi_{1}}{2\kappa}B^{\mu\nu}B_{\mu\nu}R$
$+\frac{\xi_{2}}{2\kappa}B^{\mu}{{}_{\alpha}}B^{\alpha\nu}R_{\mu\nu}+\frac{\xi_{3}}{2\kappa}B^{\mu\nu}B^{\alpha\beta}R_{\mu\nu\alpha\beta},$ (3)
where $\mathfrak{B}_{\mu\nu}=\partial_{\mu}\mathfrak{B}_{\nu}-\partial_{\nu}\mathfrak{B}_{\mu}$ and $H_{\lambda\mu\nu}=\partial_{\lambda}B_{\mu\nu}+\partial_{\mu}B_{\nu\lambda}+\partial_{\nu}B_{\lambda\mu}$ are, respectively, the field strength of the bumblebee and of the antisymmetric rank-2 tensor, $\tilde{V}(Y)$ and $V(X)$ are self-interaction potentials that drive the breaking of the Lorentz symmetry with $Y=\mathfrak{B}_{\mu}\mathfrak{B}^{\mu}\pm\mathfrak{b}^{2}$, $X=B_{\mu\nu}B^{\mu\nu}\pm b^{2}$, where $\mathfrak{b}^{2}$ and $b^{2}$ are positive real constants and the $\pm$ sign denotes whether the nature of the vector $\mathfrak{B}^{\mu}$ and of the antisymmetric tensor ${B}_{\mu\nu}$ is timelike or spacelike. Furthermore, the additional terms in the above Lagrangians account for all possible non-minimal couplings between the Lorentz-violating fields and spacetime curvature, with $R_{\alpha\beta\mu\nu}$ and $R_{\alpha\beta}$ denoting, respectively the Riemann and Ricci tensors and $\tilde{\xi}_{1,2}$ and ${\xi}_{1,2,3}$ are coupling constants. We recall that the bumblebee field does not couple to the Riemann tensor, as established in Ref. [5]. Moreover, we do not consider any coupling between the Lorentz-violating fields and the matter sector.
We assume that the self-interaction potentials $V$ and $\tilde{V}$ drive the formation of a nonzero vacuum value
$\langle\mathfrak{B}_{\mu}\rangle=\mathfrak{b}_{\mu},\quad\langle B_{\mu\nu}\rangle=b_{\mu\nu},$ (4)
which breaks local Lorentz and diffeomorphism symmetry. The minimum condition, i.e., $\langle V\rangle=0$ and $\langle\tilde{V}\rangle=0$ , is guaranteed when
$\mp\mathfrak{b}^{2}\equiv g^{\mu\nu}\mathfrak{b}_{\mu}\mathfrak{b}_{\nu},\quad\mp b^{2}\equiv g^{\mu\alpha}g^{\nu\beta}b_{\mu\nu}b_{\alpha\beta},$ (5)
where $\langle\mathfrak{B}_{\mu}\mathfrak{B}^{\mu}\rangle=\mathfrak{b}_{\mu}\mathfrak{b}^{\mu}$ and $\langle B_{\mu\nu}B^{\mu\nu}\rangle=b_{\mu\nu}b^{\mu\nu}$. It is worth emphasizing that in the vacuum state, the terms coupling the Lorentz-violating fields to the curvature scalar (On the self-consistency of compact objects in Lorentz-violating gravity theories)-(On the self-consistency of compact objects in Lorentz-violating gravity theories) can be disregarded, as they can be absorbed into the Einstein-Hilbert term through a rescaling of the gravitational constant, i.e., $\kappa\to\kappa(1\pm\tilde{\xi}_{1}\mathfrak{b}^{2}\pm\xi_{1}b^{2})^{-1}$. Since the focus of this Letter is on the solutions influenced by the anisotropies induced by the VEV of the Lorentz-violating fields, henceforth $\xi_{1}=\tilde{\xi}_{1}=0$.
By varying the action (1) with respect to the metric, one obtains the corresponding Einstein field equations, namely
$G_{\mu\nu}+\Lambda g_{\mu\nu}=\kappa T_{\mu\nu},$ (6)
where $T_{\mu\nu}=T_{\mu\nu}^{M}+T_{\mu\nu}^{LV}$ is the (total) energy-momentum tensor, with $T_{\mu\nu}^{M}$ and $T_{\mu\nu}^{LV}$ denoting the energy-momentum tensor of the matter fields and of the Lorentz-violating fields, respectively. We remark that $T_{\mu\nu}^{LV}$ is constructed by varying $S_{LV}=\int d^{4}x\sqrt{-g}(\mathcal{L}_{\mathfrak{B}}+\mathcal{L}_{{B}})$, containing, therefore, all the terms depending on $\mathfrak{B}_{\mu}$, $B_{\mu\nu}$ and their derivatives, along with all the non-minimal curvature-couplings controlled by $\tilde{\xi}_{2}$ and $\xi_{2,3}$. Additionally, by varying the action (1) with respect to the matter fields (collectively denoted by $\psi_{m}$), one obtains the equations of motion of the matter fields. Independently, by varying the action (1) with with respect to the Lorentz-violating fields, namely $\mathfrak{B}_{\mu}$ and $B_{\mu\nu}$, one obtains the field equations for the bumblebee field and for the antisymmetric rank-2 tensor field, respectively given by
$\nabla_{\alpha}\mathfrak{B}^{\alpha\nu}=$  $\,2\tilde{V}_{Y}\mathfrak{B}^{\nu}-\frac{\tilde{\xi}_{2}}{\kappa}\mathfrak{B}_{\mu}R^{\mu\nu},$ (7)
$\nabla_{\alpha}H^{\alpha\mu\nu}=$  $\,4V_{X}B^{\mu\nu}+\frac{2\xi_{2}}{\kappa}B_{\alpha}{}^{[\mu}R^{\nu]\alpha}-\frac{2\xi_{3}}{\kappa}B_{\alpha\beta}R^{\alpha\beta\mu\nu}.$ (8)
where $\tilde{V}_{Y}={\partial\tilde{V}}/{\partial Y}$ and $V_{X}={\partial V}/{\partial X}$. All these equations must be solved together in order to guarantee a self-consistent solution of the considered system.
We remark that, these consistency conditions are equivalently enforced by the conservation of the total energy-momentum tensor. By taking the covariant divergence of the Einstein equations and using the contracted Bianchi identity, $\nabla_{\mu}G^{\mu\nu}=0$, one obtains an on-shell condition that must hold:
$\nabla^{\mu}(T_{\mu\nu}^{M}+T_{\mu\nu}^{LV})=0.$ (9)
Once the equations of motion for matter and Lorentz-violating fields are satisfied, and using appropriate Bianchi identities, the above condition holds. Consequently, energy-momentum tensor conservation is guaranteed [5, 9].
Compact objects in Lorentz-violating gravity. The search for astrophysical objects in Lorentz-violating scenarios is typically based on the assumption that the Lorentz-violating fields are frozen at their VEVs. Depending on the assumption about these VEVs, different families of compact objects may arise in the Lorentz-violating scenario. It is important to point out that, imposing a particular form for the VEVs can turn the above dynamical equations into geometrical constraints, ruling out families of compact objects. These geometrical constraints emerge whenever the left-hand side of the eqs. (7)-(8) vanish. This can happen either for vanishing or (covariant) divergence-free field strength configurations. The resulting equations, for non-vanishing VEVs, imposes constraints on the Ricci and Riemann tensor.
Notably, several compact objects reported as solutions of Lorentz-violating models are based on the former assumption, namely the field strengths of the bumblebee and of the antisymmetric rank-2 tensor, in the vacuum state,—respectively $\mathfrak{b}_{\mu\nu}$ and $h_{\alpha\beta\gamma}$—, vanish. In the following we will show that this choice can, instead of allowing, forbid the emergence of some of these compact objects. We will additionally see that geometric constraints are stringent, and, for particular choices of VEV states, only a few classes of metrics are allowed.
Let us exemplify the above claim for static and spherically symmetric compact objects, that can, in general, be described by
$ds^{2}=-A(r)dt^{2}+\frac{dr^{2}}{B(r)}+r^{2}d\Omega^{2}$ (10)
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the line element of a unit sphere. If the spacetime has horizons, they are found at the zeroes of $B(r)$—for instance in black hole spacetimes. On the other hand, the zeroes of $A(r)$ are surfaces with infinite redshift. For convenience, let us consider that $B(r)=A(r)/\Omega^{2}(r)$, where $\Omega^{2}(r)$ is a positive function of $r$. We assume that the above line element satisfies the corresponding Einstein field equations of the Lorentz-violating model (1) with some matter distribution (including vacuum) that does not couple to the Lorentz-violating fields.
Among the plenty of compact objects, such as black holes and stars that can be described by the above line element, it can also represent objects endowed with a minimal hypersurface, known as wormholes. The area of the minimal hypersurface is given by $S=4\pi a^{2}$, where $a$ is the so-called throat radius. Static and spherically symmetric wormholes can be conveniently described by the Morris-Thorne line element, which is obtained by performing $\Omega^{2}(r)=A(r)/(1-s(r)/r)$, where $s(r)$ is the so-called shape function. Additionally, it is convenient to cast the redshift function as $A(r)=\exp(2\Phi(r))$. In order to describe a singularity-free and traversable wormhole, the metric function must satisfy some additional requirements. To prevent the emergence of singularities, $\Phi(r)$ must be finite everywhere, in particular when $\Phi(r)=0$ the wormhole is said to be tideless. Additionally, the throat of the wormhole is located at the minimum of the radial coordinate, $r=a$, with the shape function satisfying $s(a)=a$. Moreover, $s(r)/r<1$ for $r>a$ provides the finiteness of the proper radial distance [^1] The proper radial distance is implicitly defined by $dx/dr=\pm[1-s(r)/r]^{-1/2}$, where $\pm$ refers to the two regions connected by the wormhole throat., $x$, throughout the spacetime.
Bumblebee constraints. First, let us consider the Einstein-bumblebee model (only $\tilde{\xi}_{2}\neq 0$) and that the dynamics of the bumblebee field is frozen in its VEV state, $\langle\mathfrak{B}_{\mu}\rangle=\mathfrak{b}_{\mu}$, and additionally consider a spacelike bumblebee configurations given by
$\mathfrak{b}_{\mu}=\left(0,\mathfrak{b}\Omega(r)\sqrt{1/A(r)},0,0\right),$ (11)
such configuration was considered, for instance, in Refs. [10, 13, 16]. This configuration has vanishing field strength, i.e., $\mathfrak{b}_{\mu\nu}=0$, and therefore the equation of the bumblebee field, in the VEV state, turns into a geometrical constraint:
$2\langle\tilde{V}_{Y}\rangle\mathfrak{b}^{\nu}-\frac{\tilde{\xi}_{2}}{\kappa}\mathfrak{b}_{\mu}R^{\mu\nu}=0,$ (12)
that, specifically for this VEV configuration, in terms of the metric functions, is written as
$\Omega^{\prime}\left(rA^{\prime}+4A\right)-\Omega\left(rA^{\prime\prime}+2A^{\prime}\right)-\frac{4\kappa\langle\tilde{V}_{Y}\rangle r\Omega^{3}}{\tilde{\xi}_{2}}=0.$ (13)
The metric functions of any compact object solution of the Einstein-bumblebee model, for instance black holes, compact stars and wormholes, under the VEV configuration (11), must unavoidably satisfy the above constraint. Let us probe whether some of them are self-consistent objects within such system.
In particular, if $\Omega^{2}(r)$ is a constant, say $\Omega^{2}(r)=\omega^{2}$, the above constraint becomes
$-\omega\left(rA^{\prime\prime}+2A^{\prime}\right)-\frac{4\kappa\langle\tilde{V}_{Y}\rangle r\omega^{3}}{\tilde{\xi}_{2}}=0.$ (14)
If the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state, for instance a quadratic one, $\tilde{V}(Y)=\tfrac{1}{2}\tilde{\zeta}Y^{2}$, both $\langle\tilde{V}\rangle=0$ and $\langle\tilde{V}_{Y}\rangle=0$. Therefore, the above equation reduces to $rA^{\prime\prime}+2A^{\prime}=0$. Hence, to satisfy the constraint imposed by the bumblebee field in the considered VEV state, the redshift function must conform to
$A(r)=a_{1}-\frac{a_{2}}{r},$ (15)
where $a_{1,2}$ are integration constants that cannot be fixed solely from the constraint (12). In this scenario, only compact objects with a Schwarzschild-like redshift function are allowed by the anisotropy induced by the background bumblebee field. Particularly, one can find self-consistent black holes within the Einstein-bumblebee system. The static and spherically symmetric vacuum black hole solutions of the Einstein-bumblebee system obtained in Ref. [10] trivially satisfies the constraint (12). This black hole is described by $A(r)=1-2M/r$ and $B=A(r)/(1+l_{V})$, where $M$ is interpreted as the mass of the compact object and $l_{V}=\mathfrak{b}^{2}\tilde{\xi}_{2}$ contains the effects of the Lorentz violation.
On the other hand, if the potential is linear, say $\tilde{V}(Y)=\tfrac{1}{2}\tilde{\lambda}Y$, in the VEV state $\langle\tilde{V}\rangle$ vanishes but $\langle\tilde{V}_{Y}\rangle=\tfrac{1}{2}\lambda$. Therefore, the above equation reduces to $\tilde{\xi}_{2}(rA^{\prime\prime}+2A^{\prime})+2\kappa\omega^{2}\lambda r=0$. Consequently, to satisfy the constraint imposed by the bumblebee field in the considered VEV state, the redshift function must conform to
$A(r)=\tilde{a}_{1}-\frac{\tilde{a}_{2}}{r}-\frac{\omega^{2}\Lambda_{e}r^{2}}{3},$ (16)
that is the redshift function must have a Kottler-like form, where $\tilde{a}_{1,2}$ are integration constants that cannot be fixed solely by the constraint (12) and $\Lambda_{e}=\lambda\kappa/\tilde{\xi}_{2}$. As obtained in Ref. [13], under the same assumptions considered here, static and spherically symmetric vacuum solutions of the Einstein-bumblebee system with cosmological constant trivially satisfies the constraint (12), since the redshift function acquires the form (16), namely $A(r)=1-2M/r-(1+l_{V})\Lambda_{e}r^{2}/3$, and $B(r)=A(r)/(1+l_{V})$, where $M$ is interpreted as the mass of the compact object and $\Lambda_{e}$ an effective cosmological constant. As a consequence, black holes immersed in a universe with non-zero cosmological constant, are self-consistent objects within the Einstein-bumblebee system.
Now, let us investigate whether wormholes are self-consistent objects within the Einstein-bumblebee model under the VEV state (11). This is dictated by the following constraint
$2\tilde{\xi}_{2}r^{2}[r-s(r)][\Phi^{\prime}(r)^{2}+\Phi^{\prime\prime}(r)]+4r^{3}\kappa\langle\tilde{V}_{Y}\rangle$
$+\tilde{\xi}_{2}[s(r)-rs^{\prime}(r))][2+r\Phi^{\prime}(r)]=0,$ (17)
where we have rewritten the Eq. (13) in term of the Morris-Thorne metric functions, $\Phi(r)$ and $s(r)$. For simplicity, let us probe if tideless configurations, $\Phi(r)=0$, are self consistent in the Einstein-bumblebee model. In this case the above equation reduces to $\tilde{{\xi}_{2}}[s(r)-rs^{\prime}(r)]=-2r^{3}\kappa\langle\tilde{V}_{Y}\rangle$. It follows that, if the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state, the constraint is fulfilled only if $s(r)=s_{1}r$, where $s_{1}$ is a constant. Therefore $s(r)/r$ is constant regardless the value of $r$. It follows that, only if $s_{1}=1$ the requirement $s(a)/a=1$ holds, however this implies that $01-s(r)/r=$ 0 throughout the spacetime. If, on the contrary, $s_{1}\neq 1$ in order to satisfy $s(r)/r<1$ throughout the spacetime, $s(a)/a\neq 1$. This implies that the areal radius $r(x)$ does not have a local minimum, and therefore there is no throat in the spacetime. Consequently, if the potential that drives the breaking of the Lorentz symmetry extremizes at the vacuum state and the bumblebee field is frozen at the considered VEV configuration, tideless wormholes are not self-consistent objects in the Einstein-bumblebee model. This shows that the tideless wormhole reported in Ref. [16], although it satisfies the corresponding Einstein equations, is not self-consistent within the Einstein-bumblebee model and cannot be cast as a solution of this model.
Let us instead consider that the potential that drives the Lorentz symmetry breaking is linear, for instance $\tilde{V}(Y)=\tfrac{1}{2}\lambda Y$. In this case, the constraint is fulfilled only if $s(r)=\tilde{s}_{1}r+r^{3}\Lambda_{e}/2$, where $\tilde{s}_{1}$ is a constant and $\Lambda_{e}$ is defined above. One can check that the condition $s(a)/a=1$ holds for $\tilde{s}_{1}=1-a^{2}\Lambda_{e}/2$. Additionally, if $\Lambda_{e}<0$, the condition $01-s(r)/r>$ 0 for $r>a$ also holds. For the sake of completeness, one can compute the proper radial distance for the allowed shape function, namely $x=\pm\pi/\sqrt{2\Lambda_{e}}\pm\sqrt{2/|\Lambda_{e}|}\arctan(r/\sqrt{r^{2}-a^{2}})$, where we have set $x(a)=0$. Therefore, the square of the areal radius, in terms of the proper radial distance, reads $r^{2}(x)=a^{2}\cosh^{2}(x\sqrt{|\Lambda_{e}|/2})$, and, as one can see, has a local minimum at $x=0$. Hence, considering the linear potential, one can find self-consistent tideless wormhole configurations within the Einstein-bumblebee model. It is important to point out that such allowed wormhole configurations in the Einstein-bumblebee model do not correspond to Ellis-Bronnikov-like wormholes. This is because the areal radius of the latter relates to its proper radial distance through a parabolic equation, specifically $r^{2}(x)=a^{2}+r_{1}x^{2}$, where the constant $r_{1}=1$ in the massless Ellis-Bronnikov wormhole.
Antisymmetric rank-2 tensor constraints. Now, let us consider the Einstein-antisymmetric rank-2 tensor model ($\xi_{2}\neq 0$ and $\xi_{3}\neq 0$) and that the dynamics of the antisymmetric rank-2 tensor field is frozen in its VEV sate, $\langle B_{\mu\nu}\rangle=b_{\mu\nu}$. As demonstrated in Ref. [9], the VEV $b_{\mu\nu}$ can be decomposed as $b_{\mu\nu}=\tilde{E}_{[\mu}v_{\nu]}+\epsilon_{\mu\nu\alpha\beta}v^{\alpha}\tilde{B}^{\beta}$, where the background vectors $\tilde{E}_{\mu}$ and $\tilde{B}_{\mu}$ can be interpreted as pseudo-electric and pseudo-magnetic fields, respectively, and $v^{\mu}$ is a timelike 4-vector. The pseudo-fields $\tilde{E}_{\mu}$ and $\tilde{B}_{\mu}$ are spacelike, i.e., $\tilde{E}_{\mu}v^{\mu}=\tilde{B}_{\mu}v^{\mu}=0$.
Henceforth we consider a timelike VEV configuration given by:
$b_{\mu\nu}=\begin{pmatrix}0&-\frac{b}{\sqrt{2}}\Omega(r)&0&0\\ \frac{b}{\sqrt{2}}\Omega(r)&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},$ (18)
such configuration was considered, for instance, in Refs. [17, 18, 11, 12]. With this choice, the vacuum condition is satisfied as expected, and the field strength vanishes, i.e., $h_{\mu\nu_{\lambda}}=0$. Thus, the equation of the antisymmetric rank-2 tensor field, in the VEV state, turns into the following geometric constraint
$4\langle V_{X}\rangle b^{\mu\nu}+\frac{2\xi_{2}}{\kappa}b_{\alpha}{}^{[\mu}R^{\nu]\alpha}-\frac{2\xi_{3}}{\kappa}b_{\alpha\beta}R^{\alpha\beta\mu\nu}=0,$ (19)
that, in terms of the metric functions, specifically for the VEV configuration (18), is written as
$\xi_{2}\left(r\Omega A^{\prime\prime}-rA^{\prime}\Omega^{\prime}+2\Omega A^{\prime}-2A\Omega^{\prime}\right)$ (20)
$+\xi_{3}\left(2r\Omega A^{\prime\prime}-2rA^{\prime}\Omega^{\prime}\right)+4\kappa r\langle V_{X}\rangle\Omega^{3}=0.$
This constraint must be obeyed for any compact object solution of the Einstein-antisymmetric rank-2 tensor model, for instance black holes, compact stars and wormholes, under the VEV configuration (18).
Particularly, if $\Omega^{2}(r)$ is a constant, for instance $\Omega^{2}(r)=\omega^{2}$, the constraint (20) reduces to
$rA^{\prime\prime}(\xi_{2}+2\xi_{3})+2\xi_{2}A^{\prime}+4\kappa r\langle V_{X}\rangle\omega^{2}=0.$ (21)
By considering that the potential that drives the breaking of the Lorentz symmetry extremizes at the VEV state, for example a quadratic one, both $\langle{V}\rangle$ and $\langle{V}_{X}\rangle$ vanish. Thus, the above equation becomes $rA^{\prime\prime}(\xi_{2}+2\xi_{3})+2\xi_{2}A^{\prime}=0$. Consequently, to satisfy the constraint imposed by $b_{\mu\nu}$, the redshift function must conform to
$A(r)=c_{1}+\frac{c_{2}r^{\gamma}}{\gamma},$ (22)
where $\gamma=1-{2\xi_{2}}/(\xi_{2}+2\xi_{3})$ and ${c}_{1,2}$ are integration constants that cannot be fixed solely from the constraint (8). It follows that depending on the combination of the coupling constants, different families of functions $A(r)$ are allowed by the model. For instance, if $\xi_{3}=0$, $\gamma=-1$, and consequently $A(r)=c_{1}-c_{2}/r$, which has a Schwarzschild-like form. On the other hand, if $\xi_{2}=0$, $\gamma=1$, and consequently $A(r)=c_{1}+c_{2}r$, where the linear term can be understood as a sort of Rindler-type acceleration [19, 20]. Since $\gamma$ can take arbitrary real values, even an (A)dS-like redshift function is allowed, for instance if $\xi_{3}=-3\xi_{2}/2$, $\gamma=2$, and $A=c_{1}+c_{2}r^{2}/2$. Although redshift functions with arbitrary exponents can indeed satisfy the above constraint, finding reasonable matter configurations that give support to such configurations is difficult, therefore they have limited physical relevance.
We emphasize that black hole solutions in the Einstein-antisymmetric rank-2 tensor system—under the VEV state considered here—have already been reported in the literature. Specifically, the black hole in Ref. [11], defined by the metric functions $A(r)={1}/{(1-l_{T})}-{2M}/{r}$ and $B(r)=A(r)$, and the solution in Ref. [12], described by $A(r)=1-{2M}/{r}$ and $B(r)={A(r)}/{(1-l_{T})}$, both conform to the constraint (21). Here, $l_{T}=b^{2}\tilde{\xi}_{2}/2$ encodes Lorentz-violating effects, while $M$ represents the mass of the compact object. Therefore, these black holes are self-consistent within the Einstein-antisymmetric rank-2 tensor model. Conversely, in Ref. [17], it was reported a vacuum black hole solution, described by $A(r)=1-2M/r+\Upsilon/r^{2/l_{T}}$ and $B(r)=A(r)$, with $\Upsilon$ being a constant, that, as we can see from Eq. (22), does not conform to the constraint (21). Hence, this black hole is not self-consistent within the considered model and therefore cannot be cast as a solution of it.
Assuming, instead, a linear self-interaction potential, at the VEV state, $\langle V_{X}\rangle=\frac{1}{2}\lambda$, one finds that the redshift function must conform to
$A(r)=\tilde{c}_{1}+\frac{\tilde{c}_{2}r^{\gamma}}{\gamma}-\frac{\omega^{2}\tilde{\Lambda}_{e}}{3}r^{2}$ (23)
in order to respect the anisotropies induced by $b_{\mu\nu}$. Here $\tilde{c}_{1,2}$ are integration constants that cannot be fixed solely from the constraint (8) and $\tilde{\Lambda}_{e}=\frac{\kappa\lambda}{3\xi_{2}+{2}\xi_{3}/{3}}$, with $\xi_{3}\neq-3\xi_{2}/2$. Particularly interesting cases cases are given by $\xi_{3}=0$, that leads to $\gamma=-1$, $\tilde{\Lambda}_{e}=\kappa\lambda/2\xi_{2}$ and a a Kottler-like redshift function, and $\xi_{2}=0$, that leads to $\gamma=1$, $\tilde{\Lambda}_{e}=\kappa\lambda/3\xi_{3}$ and a redshift function with an effective cosmological constant and a Rindler-type acceleration, respectively. We point out that—under the VEV state considered here—static and spherically symmetric vacuum solutions of the Einstein-antisymmetric rank-2 tensor system with cosmological constant have been reported in the literature [11, 12]. We remark that these solutions satisfy the constraint (23).
Now, let us investigate whether wormholes are self-consistent objects within curvature-coupled antisymmetric rank-2 tensor models under the VEV state (18). This is achieved if the wormhole metric functions satisfy the constraint
$\xi_{3}\{2r[s(r)+rs^{\prime}(r)]\Phi^{\prime}(r)+4r^{2}[r-s^{\prime}(r)][\Phi^{\prime}(r)^{2}-\Phi^{\prime\prime}(r)]\}$
$+{\xi}_{2}\{s(r)-rs^{\prime}(r)-[s(r)+r(s^{\prime}(r)-2)]r\Phi^{\prime}(r)\}$
$0\displaystyle 2{\xi}_{2}r^{2}[r-s(r)][\Phi^{\prime}(r)^{2}+\Phi^{\prime\prime}(r)]+4r^{3}\kappa\langle{V}_{X}\rangle=$ 0 . (24)
In particular, let us probe if tideless wormholes are self-consistent objects in such a system. By considering $\Phi(r)=0$, the above constraint simplifies to $\xi_{2}[s(r)-rs^{\prime}(r)]=-4r^{3}\kappa\langle V_{X}\rangle$. Therefore, considering that the potential extremizes at the VEV state, one finds that tideless wormholes are not self-consistent objects within the Einstein-antisymmetric rank-2 tensor model, once the areal radius $r(x)$ cannot have a local minimum, similarly as the constraint present in the Einstein-bumblebee model. However, by considering a linear potential, $V(X)=\frac{1}{2}\lambda X$, one is able to find self-consistent tideless wormholes. Specifically, their shape function must behave as $s(r)=(1-a^{2}\tilde{\Lambda}_{e})r+\tilde{\Lambda}_{e}r^{3}$. In particular, our results show that the tideless wormhole reported in Ref. [18], although it satisfies the corresponding Einstein equations, is not self-consistent within the Lorentz-violating curvature-coupled antisymmetric rank-2 tensor model and cannot be cast as a solution of this model.
Remarkably, the whole term multiplying $\xi_{3}$ in Eq. (On the self-consistency of compact objects in Lorentz-violating gravity theories) vanishes for tideless configurations, regardless of the form of the shape function. It implies that, in Riemann-coupled antisymmetric rank-2 tensor models ($\xi_{2}=0$ and $\xi_{3}\neq 0$), considering self-interaction potential that extremizes at the VEV state, the constraint (On the self-consistency of compact objects in Lorentz-violating gravity theories) is identically satisfied. However, if instead a linear self-interaction potential is considered, tideless wormholes are ruled out by the above constraint. Notably, a tideless wormhole of this system was reported in Ref. [15]. The authors find a tideless wormhole configuration sourced by an ideal fluid, considering a potential that extremizes at the VEV state, which by the above analysis is a self-consistent object of the Riemann-coupled antisymmetric rank-2 tensor model.
We remark that the solution reported in Ref. [15] is not the unique possible tideless wormhole supported within this framework, and other tideless wormholes, associated to different shape functions, can be found. Surprisingly, a Lorentz-violating version of the (massless) Ellis-Bronnikov wormhole, namely a tideless wormhole with shape function $s(r)\propto a^{2}/r$, is still missing in the literature of Lorentz-violating bumbeblee or antisymmetric rank-2 tensors fields, and due to our analysis, such wormhole is self consistent within the latter Lorentz-violating gravity framework. A derivation of such wormhole and other viable wormholes that satisfy the above discussed criteria is currently underway.
Final words. The self-consistency of compact objects in Lorentz-violating gravity theories demands not only the fulfillment of Einstein and matter field equations but also the satisfaction of the Lorentz-violating field equations. These equations can impose stringent geometric constraints in vacuum states, which we argue may rule out the existence of entire families of compact object solutions in such frameworks. By analyzing two prominent models—spontaneous Lorentz symmetry breaking via a bumblebee vector field and an antisymmetric rank-2 tensor field—we demonstrate that common assumptions, such as freezing Lorentz-violating fields at their VEVs while vanishing their field strength, may obstruct rather than enable the emergence of certain compact objects. Our results challenge the physical admissibility of several previously reported solutions, including black holes and wormholes, as their failure to satisfy the Lorentz-violating field equations leads to inconsistencies such as non-conservation of the total energy-momentum tensor.
A promising avenue to circumvent these constraints lies in coupling Lorentz-violating fields to matter. Such couplings can modify the dynamics of these fields in their vacuum state, potentially permitting a broader class of solutions through induced anisotropies. For instance, coupling a $U(1)$ gauge field to Lorentz-violating fields has successfully generated self-consistent Lorentz-violating analogues of Reissner-Nordström black holes [21, 22]. This approach opens a pathway to extend well-known GR solutions into Lorentz-violating frameworks while preserving self-consistency. In particular, Lorentz-violating analogues of the Ellis-Bronnikov wormholes, excluded in the Einstein-bumblebee model considered here under vacuum conditions, could become viable with appropriate matter couplings. A detailed investigation of such wormhole solutions in the Einstein-bumblebee model is in progress, with results to be reported in due course.
Acknowledgements.
The authors would like to acknowledge Fundação de Amparo à Pesquisa e ao Desenvolvimento Científico e Tecnológico do Maranhão (FAPEMA), Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) – Finance Code 001, from Brazil, for partial financial support. L.A.L is supported by FAPEMA BPD- 08975/24.